OpenStudy (anonymous):

This question... help me and you'll be rewarded with a metal and a side of fries

OpenStudy (anonymous):

Brian is creating a collage on a piece of cardboard that has an area of 120r3 square centimeters. The collage is covered entirely by pieces of paper that do not overlap. Each piece has an area of sqrt(r^5) square centimeters. Use the given information to determine an expression for the total number of pieces of paper used.

OpenStudy (anonymous):

i think you need to compute \[120 r^3\div \sqrt{r^5}\]

OpenStudy (anonymous):

or \[\frac{120r^3}{\sqrt{r^5}}\]

OpenStudy (anonymous):

if i read it correctly start with \[\sqrt{r^5}=r^2\sqrt{r}\]

OpenStudy (anonymous):

they really reach for these word problems, don't they?

OpenStudy (anonymous):

can i have pepper on the fries?

OpenStudy (anonymous):

Yes, thank you! I understand it. It looks like I was right up until dividing which I have to remember :)

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

WAIT. Don't forget your fries!

OpenStudy (anonymous):

ooh pepper and mayo!!

OpenStudy (anonymous):

;)

OpenStudy (anonymous):

oh crap, i have to give the fries back, i made a mistake! there is no 2 in the denominator, so it is just \[120r\sqrt{r}\]

OpenStudy (anonymous):

let me try this again

OpenStudy (anonymous):

\[\frac{120r^3}{\sqrt{r^5}}=\frac{120r^3}{r^2\sqrt{r}}\] \[=\frac{120r}{\sqrt{r}}=120\sqrt{r}\] that's better!

OpenStudy (anonymous):

i have no idea where i got the 2 from, it was just a dumb mistake on my part

OpenStudy (anonymous):

Its fine, thank you and at least you caught it :)

OpenStudy (anonymous):

pretty sure it is right now i hope you didn't submit the wrong thing on line or something